# Matching-Based Policy Learning

### v1

@@ A.7.4 Step 4: High probability bound for $A(\pi^{*})-A(\hat{\pi})$
Combining (14), (15), (A.7.3), (17), (18), (21) and (33), it follows that for any $\delta>0$, with at least $1-9\delta$ and for all $n>\max\{N(c_{3}),N(c_{4},\delta),N(c_{6},q)\}$, we have
@@
Note that $\frac{1}{n}\sqrt{\sum_{i=1}^{n}K^{2}_{M}(i)}=\frac{1}{\sqrt{n}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}K^{2}_{M}(i)}=O_{p}\left(\frac{1}{\sqrt{n}}\right)=o_{p}\left(\frac{1}{\sqrt{n^{1-1/q}}}\right)$, where the second equality follows from Markov’s inequality. Let $c_{7}>0$, then for any $\delta>0$, there exists $N(c_{7},\delta)$ such that for all $n>N(c_{7},\delta)$, we have
@@
Similarly, since $\frac{1}{n}\max_{i\leq n}K_{M}(i)=o_{p}\left(\frac{1}{n^{1-1/q}}\right)$ by Markov’s inequality. Let $c_{8}>0$, there exists $N(c_{8},\delta)$ such that for all $n>N(c_{8},\delta)$, we have
@@
Recall that $A(\pi^{*})-A(\hat{\pi})=2R(\hat{\pi})$. Let $N_{\delta,q}=\max\{N(c_{3}),N(c_{4},\delta),N(c_{6},q),N(c_{7},\delta),N(c_{8},\delta)\}$, it follows from (34), (35), and (36) that, for any $\delta>0$ and any $q>1$, with at least probability $1-11\delta$ and for all $n>N_{\delta,q}$, we have
@@ B.1 Data Generating Process
Table S1 presents detailed information for the data generating processes in simulation studies. Additionally, the proportions of the controlled versus treated units are displayed. For the propensity score model, Scenario 1 is the same as Kang and Schafer (2007), where $l(\cdot)$ takes a linear form relative to its argument and the proportion is balanced. In Scenario 2 we consider a nonlinear form of $l(\cdot)$, while the proportion is still balanced. Scenario 3 is modified from Scenario 1 to introduce imbalanced proportion. We consider a randomized experiment in Scenario 4, where the proportion is extremely imbalanced. Scenario 5 is modified from Wu et al. (2020), with a nonlinear propensity score model and extremely imbalanced proportion.
Table S1: Data generating processes for different scenarios in simulation studies. Outcome variable $Y=m(X)+Wc(X)+Z$, where binary treatment variable $W$ generated by $\Pr(W=1|X)=\exp\left(l(X)\right)/\left[1+\exp\left(l(X)\right)\right]$.
@@ B.2 Simulation Results with Nonlinear Main Effect
B.3 Comparison with Policy Tree
@@ Appendix C Additional Results in Real Data Application
Figure S5: Optimal policies estimated by using all the samples in DW dataset. (a) The policy derived from MB-LASSO-M5 depends only on pre-treatment earnings, and 401 (90%) of the samples are predicted to have "exposed" as the optimal choice. (b) The policy derived from Policy Tree depends on both pre-treatment earnings and education, and 290 (65%) of the samples are predicted to have "exposed" as the optimal choice.
@@
Figure S6: Optimal policies estimated by using all the samples in DW-CPS3 dataset. (a) The policy derived from MB-M5 depends on age, education and pre-treatment earnings, and 560 (64%) of the samples are predicted to have "exposed" as the optimal choice. (b) The policy derived from Policy Tree depends on age and pre-treatment earnings, and 557 (64%) of the samples are predicted to have "exposed" as the optimal choice.

### v2

@@ A.7.4 Step 4: High probability bound for $A(\pi^{*})-A(\hat{\pi})$
Combining (15), (16), (A.7.3), (18), (19), (22) and (34), it follows that for any $\delta>0$, with probability at least $1-9\delta$ and for all $n>\max\{N(c_{3}),N(c_{4},\delta),N(c_{6},q)\}$, we have
@@
Note that $\frac{1}{n}\sqrt{\sum_{i=1}^{n}K^{2}_{M}(i)}=\frac{1}{\sqrt{n}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}K^{2}_{M}(i)}=O_{p}\left(\frac{1}{\sqrt{n}}\right)=o_{p}\left(\frac{1}{\sqrt{n^{1-1/q}}}\right)$, where the second equality follows from Markov’s inequality. For $c_{7}>0$, then for any $\delta>0$, there exists $N(c_{7},\delta)$ such that for all $n>N(c_{7},\delta)$, we have
@@
Similarly, since $\frac{1}{n}\max_{i\leq n}K_{M}(i)=o_{p}\left(\frac{1}{n^{1-1/q}}\right)$ by Markov’s inequality. For $c_{8}>0$, there exists $N(c_{8},\delta)$ such that for all $n>N(c_{8},\delta)$, we have
@@
Recall that $A(\pi^{*})-A(\hat{\pi})=2R(\hat{\pi})$. Let $N_{\delta,q}=\max\{N(c_{3}),N(c_{4},\delta),N(c_{6},q),N(c_{7},\delta),N(c_{8},\delta)\}$. It follows from (35), (36), and (37) that, for any $\delta>0$, $q>1$ and $n>N_{\delta,q}$, with probability at least $1-11\delta$, the following inequality holds:
@@ B.1 Data Generating Process
Table S1 presents detailed information for the data generating processes in simulation studies. Additionally, the proportions of the controlled versus treated units are displayed. For the propensity score model, Scenario 1 is the same as Kang and Schafer (2007), where $l(\cdot)$ is a linear combination of predictors and the proportion is balanced. In Scenario 2 we consider a nonlinear form of $l(\cdot)$, while the proportion is still balanced. Scenario 3 is modified from Scenario 1 to introduce imbalanced proportion. We consider a randomized experiment in Scenario 4, where the proportion is extremely imbalanced. Scenario 5 is modified from Wu et al. (2020), where a nonlinear propensity score model is in use and the proportions in the two groups are extremely imbalanced.
Table S1: Data generating processes for different scenarios in simulation studies. The outcome model is $Y=m(X)+Wc(X)+e$, where the binary treatment variable $W$ is generated by $\Pr(W=1|X)=\exp\left(l(X)\right)/\left[1+\exp\left(l(X)\right)\right]$.
@@ B.2 Simulation Results with Nonlinear Main Effect
B.3 Comparison of MB-Learning and Policy Tree with Sample Size Varying from 200 to 8000
@@ Appendix C Additional Results in Real Data Application
Figure S5: Optimal policies estimated by using all the data in DW dataset. (a) The optimal policy derived from MB-LASSO-M5 depends only on pre-treatment earnings, and 401 (90%) of the job-seekers are recommended to be "exposed". (b) The optimal policy derived from Policy Tree depends on both pre-treatment earnings and education, and 290 (65%) of the job-seekers are recommended to be "exposed".
@@
Figure S6: Optimal policies estimated by using all the data in DW-CPS3 dataset. (a) The optimal policy derived from MB-M5 depends on age, education and pre-treatment earnings, and 560 (64%) of the samples are recommended to be "exposed". (b) The optimal policy derived from Policy Tree depends on age and pre-treatment earnings, and 557 (64%) of the samples are recommended to be "exposed".